# An efficient Monte Carlo method for valid prior-free possibilistic statistical inference

### v2

@@ Abstract
Inferential models (IMs) offer prior-free, Bayesian-like posterior degrees of belief designed for statistical inference, which feature a frequentist-like calibration property that ensures reliability of said inferences. The catch is that IMs’ degrees of belief are possibilistic rather than probabilistic and, since the familiar Monte Carlo methods approximate probabilistic quantities, there are computational challenges associated with putting this framework into practice. The present paper addresses these challenges by developing a new Monte Carlo method designed specifically to approximate the IM’s possibilistic output. The proposal is based on a characterization of the possibilistic IM’s credal set, which identifies the “best probabilistic approximation” of the IM as a mixture distribution that can be readily approximated and sampled from. These samples can then be transformed into an approximation of the possibilistic IM. Numerical results are presented highlighting the proposed approximation’s accuracy and computational efficiency.
Keywords and phrases: confidence distribution; credal set; Gaussian possibility; inferential model; variational inference.
@@ 1 Introduction
Numerous efforts to advance Fisher’s vision of fiducial inference have been made over the years. Among these efforts, arguably “one of the original statistical innovations of the 2010s” (Cui and Hannig, 2024) is the inferential model (IM) framework put forward in Martin and Liu, (2013); Martin and Liu, 2015a ; Martin and Liu, 2015c and later synthesized in Martin and Liu, 2015b . What’s unique about the IM framework is that its output takes the form of an imprecise probability—specifically, a possibility measure—and, consequently, inference is based on possibilistic reasoning: hypotheses assigned small possibility are refuted by the data while hypotheses whose complements are assigned small possibility are corroborated by the data. This shift from probability to possibility theory has principled motivations: the IM’s possibilities satisfy a frequentist-style validity property that Bayesian/fiducial probabilities do not satisfy, i.e., the IM’s possibility assigned to true hypotheses tends to be not small, hence it’s a provably rare event that the data-driven IM output refutes a true hypothesis or corroborates a false hypothesis. More details about IMs and their properties are provided in Section 2 below. Despite all that the IM framework has to offer, it has been slow to gain traction in the statistical community, largely due to computational challenges. Indeed, as shown recently in Jacob et al., (2021), the familiar Monte Carlo methods used to approximate ordinary probabilities simply aren’t enough to approximate imprecise probabilities; something more is needed. The goal of this paper is to identify the aforementioned “something more” for the case of possibilistic IMs and to develop the corresponding Monte Carlo machinery needed to make these computations, and the IM framework more generally, readily available for everyday use in applications.
@@
The jumping off point here is the fact that all (coherent) imprecise probabilities, including the possibilistic IM’s output, correspond to a (non-empty, closed, and convex) set of ordinary or precise probabilities, called the credal set. Different brands of imprecise probabilities have their own mathematical properties, and one way this distinction manifests is in the kinds of constraints imposed on their associated credal sets. Possibility measures are among the simplest imprecise probabilities and, in turn, their credal sets have a relatively simple characterization. In particular, results like that in Theorem 1 below make it relatively easy to identify probability distributions in the credal set that “best approximate” the IM’s possibilistic output. The best approximation is relatively simple mathematically but might still be difficult to compute; fortunately, further approximations can be made. Such second-level approximations can take various forms but, inspired by the asymptotic results in Martin and Williams, (2024), here I’ll use a family of Gaussian variational approximations as proposed in Cella and Martin, (2024). These variational approximations on their own are not fully satisfactory, but here I show they can be stitched together in a satisfactorily way. Specifically, my proposal is to follow the guidelines suggested in the characterization result in Theorem 1 to construct a probability distribution, with certain Gaussian-like features, that is no more concentrated than this best probabilistic approximation of the IM’s possibilistic output. The aforementioned Gaussian-like features make this easy and fast to sample from, thereby facilitating principled probabilistic approximations to the IM’s output. I also show how this probabilistic approximation can be readily transformed into a possibility measure that is both easy to compute and accurately approximates the target IM output.
@@
The remainder of this paper is organized as follows. Section 2 provides the necessary background on possibilistic IMs, highlighting the existing computational bottlenecks. The credal set characterization and proposed Monte Carlo strategy are presented in Section 3. Three illustrative examples, including logistic regression, are presented in Section 4. These are cases where direct-but-naive IM computations are expensive but not out of reach, so the accuracy of the proposed approximation can be assessed. As a more challenging application, Section 5 presents a non-trivial analysis of a small, censored data set involving a semiparametric model. For this application, the direct-but-naive IM solution is too expensive for practitioners to take seriously but the proposed Monte Carlo-based solution can be carried out in a matter of seconds. The insights and results presented here make up an important first step in a series of developments needed to create an IM toolbox for practitioners. Among other things, the concluding remarks in Section 6 describe my vision of where these developments can and need to go. Some relevant technical details are given in an appendix, including a description of the variational approximation developed in Cella and Martin, (2024).
@@ 2 Background
$R(z,\theta)=\frac{L_{z}(\theta)}{\sup_{\vartheta}L_{z}(\vartheta)}.$
I will implicitly assume here that the denominator is finite for almost all $z$. As is typical in the literature, I will also assume that prior information about $\Theta$ is vacuous.
The relative likelihood itself defines a data-dependent possibility contour, i.e., it is a non-negative function such that $\sup_{\theta}R(z,\theta)=1$ for almost all $z$. This contour, in turn, determines a possibility measure that can be used for uncertainty quantification about $\Theta$, given $Z=z$, which has been extensively studied in the literature (e.g., Shafer, 1982; Wasserman, 1990a ; Denœux, 2006, 2014). In particular, these references suggest assigning data-driven possibility values to hypotheses $H$ about $\Theta$ via the rule
@@
$\overline{\Pi}_{z}(H)=\sup_{\theta\in H}\pi_{z}(\theta),\quad H\subseteq\mathbb{T}.$ (2)
It won’t be needed in the present paper, but there’s a corresponding necessity measure, or lower probability, defined via conjugacy: $\underline{\Pi}_{z}(H)=1-\overline{\Pi}_{z}(H^{c})$. An essential feature of this IM construction is its so-called validity property:
@@
$\sup_{\Theta\in H}\mathsf{P}_{\Theta}\bigl{\{}\overline{\Pi}_{Z}(H)\leq\alpha\bigr{\}}\leq\alpha,\quad\text{all $\alpha\in[0,1]$, all $H\subseteq\mathbb{T}$}.$ (5)
In words, a valid IM assigns possibility $\leq\alpha$ to true hypotheses at rate $\leq\alpha$ as a function of data $Z$. This gives the IM its “inferential weight”—(5) implies that $\overline{\Pi}_{z}(H)$ is not expected to be small when $H$ is true, so one is inclined to doubt the truthfulness of a hypothesis $H$ if $\overline{\Pi}_{z}(H)$ is small. Third, the above property ensures that the possibilistic IM is safe from false confidence (Balch et al., 2019; Martin, 2019; Martin, 2024b ), unlike all default-prior Bayes and fiducial solutions. An even stronger, uniform-in-hypotheses version of (5) holds, which offers opportunities for the data analyst to do more with the IM than test prespecified hypotheses (Cella and Martin, 2023). For further details about possibilistic IMs’ properties, including de Finetti-style no-sure-loss properties, its connection to Bayesian/fiducial inference, etc, see Martin, 2022b ; Martin, 2023b ; Martin, 2023a .
In a Bayesian analysis, inference is based on summaries of the posterior distribution, e.g., posterior probabilities of scientifically relevant hypotheses, expectations of loss/utility functions, etc. All of these summaries boil down to integration involving the probability density function that determines the posterior. Virtually the same statement holds for the possibilistic IM: the lower–upper probability pairs for scientifically relevant hypotheses, lower–upper expectations of loss/utility functions (Martin, 2021), etc. involve optimization of the possibility contour $\pi_{z}$. More precisely, the proper calculus for possibilistic reasoning is Choquet integration (e.g. Troffaes and de Cooman, 2014, App. C), which goes as follows. If $h:\mathbb{T}\to\mathbb{R}$ is a non-negative function, then the Choquet integral with respect to the possibility measure $\overline{\Pi}_{z}$ is defined as
[…]
As an important special case, suppose that $h(\theta)=1(\theta\in H)$, where $H=\{\theta:k(\theta)\in K\}$ for some function $k$ and some subset $K\subseteq k(\mathbb{T})$. Then
$\overline{\Pi}_{z}h=\sup_{\theta:h(\theta)=1}\pi_{z}(\theta)=\sup_{\theta\in H}\pi_{z}(\theta)=\sup_{\kappa\in K}\sup_{\theta:k(\theta)=\kappa}\pi_{z}(\theta).$
There’s a clear analogy that can be made between the right-hand side of the above display and the more familiar Bayesian calculus: the degree of possibility/probability assigned to “$\Theta\in H$” or, equivalently, to “$k(\Theta)\in K$” is obtained by first getting the marginal contour/density for $k(\Theta)$ at $\kappa$ by optimizing/integrating over $\{\theta:k(\theta)=\kappa\}$ and then optimizing/integrating over $\kappa\in K$. Unlike with Bayesian integration, however, the IM’s optimization operation ensures that the validity property inherent in $\pi_{z}$ is transferred to the marginal IM for $k(\Theta)$, which implies that it’s safe from false confidence.
@@
where $Z_{m,\theta}$ are independent copies of the data $Z$, specifically drawn from $\mathsf{P}_{\theta}$, for $m=1,\ldots,M$. The above computation is feasible at one or a few different $\theta$ values, but this often needs to be carried out over a fine grid covering the relevant portion of the parameter space $\mathbb{T}$. For example, identifying the confidence set in (4) requires finding which $\theta$’s satisfy $\pi_{z}(\theta)\geq\alpha$, and a naive approach is to compute the contour over a huge grid and then keep those that (approximately) meet the aforementioned condition. This amounts to lots of wasted and expensive computations. More generally, the relevant summaries of the IM output involve optimization, as in (6), and doing so numerically requires many contour function evaluations. This is a serious bottleneck, so new and not-so-naive computational strategies are desperately needed.
@@ 3 Monte Carlo methods for IMs
3.1 Key insights
Given a possibilistic IM $z\mapsto(\underline{\Pi}_{z},\overline{\Pi}_{z})$ with contour function $\pi_{z}$, I’ll refer to the level sets $C_{\alpha}(z)=\{\theta\in\mathbb{T}:\pi_{z}(\theta)\geq\alpha\}$ defined in (4) as $\alpha$-cuts. According to the theory reviewed in Section 2, the $\alpha$-cut $C_{\alpha}(z)$ is a nominal $100(1-\alpha)$confidence region for $\Theta$. Aside from the practical utility of confidence regions, there’s much more information in the IM’s $\alpha$-cuts. Towards this, the credal set $\mathscr{C}(\overline{\Pi}_{z})$ corresponding to the IM’s $\overline{\Pi}_{z}$ is the set of (possibly data-dependent) probability measures that it dominates, i.e.,
$\mathscr{C}(\overline{\Pi}_{z})=\{\mathsf{Q}_{z}\in\text{probs}(\mathbb{T}):\text{$\mathsf{Q}_{z}(H)\leq\overline{\Pi}_{z}(H)$ for all measurable $H$}\},$
where $\text{probs}(\mathbb{T})$ is the set of probability measures on $\mathbb{T}$. Then the claim is that the $\alpha$-cuts determine the IM’s credal set. Indeed, there’s the following well-known characterization (e.g., Couso et al., 2001; Destercke and Dubois, 2014) of the credal set $\mathscr{C}(\overline{\Pi}_{z})$:
$\mathsf{Q}_{z}\in\mathscr{C}(\overline{\Pi}_{z})\iff\mathsf{Q}_{z}\{C_{\alpha}(z)\}\geq 1-\alpha\quad\text{for all $\alpha\in[0,1]$}.$ (8)
That is, a probability distribution $\mathsf{Q}_{z}$ belongs to the IM’s credal set if and only if it’s a confidence distribution, i.e., it assigns probability $\geq 1-\alpha$ to each of the IM’s $\alpha$-cuts. More about this notion of “confidence distributions” and when Bayesian/fiducial distributions achieve equality in (8) can be found in Martin, 2023a . At a soon-to-be-demonstrated practical level, elements of the IM’s credal set have a relatively simple characterization, as the following theorem demonstrates. The result is presented in a more general, not-IM-specific context because this simplifies the notation.
@@ 3.1 Key insights
Let $\overline{\Pi}$ be a possibility measure on a space $\mathbb{T}$, with corresponding contour $\pi$ and $\alpha$-cuts $C_{\alpha}=\{\theta\in\mathbb{T}:\pi(\theta)\geq\alpha\}$, $\alpha\in(0,1)$. Then $\mathsf{Q}\in\mathscr{C}(\overline{\Pi})$ if and only if
@@
See Appendix A.1. ∎
I would be surprised if the characterization in Theorem 1 is genuinely new, but I’ve not found a reference for this result exactly. A similar result is presented in Wasserman, 1990b (, Theorem 2.1) characterizing the credal set of a belief function determined by a given distribution and set-valued mapping. Hose, (2022, Eq. 2.44) also gives a special case of the right-hand side of (9) but makes no claims that his version is a complete characterization of $\mathscr{C}(\overline{\Pi})$; what I specifically propose to do in (12) below closely matches the Hose’s formula. As an aside, choosing the kernel $\mathsf{K}^{\beta}$ to be a uniform distribution supported on $C_{\beta}$ is an idea that commonly appears in the literature, e.g., the pignistic probability (Smets and Kennes, 1994) or the Shapley value (Shapley, 1953) of a game.
The above theorem describes the contents of the credal set $\mathscr{C}(\overline{\Pi})$ corresponding to a possibility measure $\overline{\Pi}$. A relevant follow-up question is if there exists an element in $\mathscr{C}(\overline{\Pi})$ that’s “maximally consistent” with $\overline{\Pi}$, i.e., a distribution $\mathsf{Q}^{\star}$ such that
@@
Such a probability distribution will be called an inner probabilistic approximation of $\overline{\Pi}$. The theorem’s proof offers some insights that help to answer this question. Indeed, the two properties needed to achieve the equality in (10) are, first, that the kernel satisfies $\mathsf{K}^{\beta}(C_{\alpha})=0$ for all pairs $(\alpha,\beta)$ with $\beta<\alpha$ and, second, that $\mathsf{M}={\sf Unif}(0,1)$.
As an illustration, which will prove to be useful in Section 3.3 below, consider the Gaussian possibility measure $\overline{\Pi}$, with corresponding mean vector $m$ and covariance matrix $V$, defined in Martin and Williams, (2024) with contour
@@
where $F_{d}$ is the ${\sf ChiSq}(d)$ distribution function. Then the corresponding $\alpha$-cuts are given by the $d$-dimensional ellipsoids

### v3

@@ Abstract
Inferential models (IMs) offer prior-free, Bayesian-like posterior degrees of belief designed for statistical inference, which feature a frequentist-like calibration property that ensures reliability of said inferences. The catch is that IMs’ degrees of belief are possibilistic rather than probabilistic and, since the familiar Monte Carlo methods approximate probabilistic quantities, there are significant computational challenges associated with putting this framework into practice. The present paper overcomes these challenges by developing a new Monte Carlo method designed specifically to approximate the IM’s possibilistic output. The proposal is based on a characterization of the possibilistic IM’s credal set, which identifies the “best probabilistic approximation” of the IM as a mixture distribution that can be readily approximated and sampled from. These samples can then be transformed into an approximation of the possibilistic IM. Numerical results are presented highlighting the proposed approximation’s accuracy and computational efficiency.
Keywords and phrases: confidence distribution; credal set; Gaussian possibility; inferential model; inner probabilistic approximation.
@@ 1 Introduction
Numerous efforts to advance Fisher’s vision of fiducial inference have been made over the years. Among these efforts, arguably “one of the original statistical innovations of the 2010s” (Cui and Hannig, 2024) is the inferential model (IM) framework put forward in Martin and Liu, (2013); Martin and Liu, 2015a ; Martin and Liu, 2015c and later synthesized in Martin and Liu, 2015b . What’s unique about the IM framework is that its output takes the form of an imprecise probability—specifically, a possibility measure—and, consequently, inference is based on possibilistic reasoning: hypotheses assigned small possibility are refuted by the data while hypotheses whose complements are assigned small possibility are corroborated by the data. This shift from probability to possibility theory has principled motivations: the IM’s possibilities satisfy a frequentist-style validity property that Bayesian/fiducial probabilities do not satisfy, i.e., the IM’s possibility assigned to true hypotheses tends to be not small, hence it’s a provably rare event that the data-driven IM output refutes a true hypothesis or corroborates a false hypothesis. More details about IMs and their properties are provided in Section 2 below; some basic background on possibility theory is presented in Appendix A. Despite all that the IM framework has to offer, it has been slow to gain traction, largely due to computational challenges. Indeed, as shown recently in Jacob et al., (2021), the familiar Monte Carlo methods used to approximate ordinary probabilities aren’t enough to approximate imprecise probabilities; something more is needed. The goal of this paper is to identify the aforementioned “something more” for possibilistic IMs and to develop the corresponding computational machinery needed to make the IM framework accessible for everyday use in applications.
@@
The jumping off point here is the fact that all (coherent) imprecise probabilities, including the possibilistic IM’s output, correspond to a (non-empty) set of ordinary or precise probabilities, called the credal set. Different brands of imprecise probabilities have their own mathematical properties, and one way this distinction manifests is in the kinds of constraints imposed on their associated credal sets. Possibility measures are among the simplest imprecise probabilities and, in turn, their credal sets have a relatively simple characterization. In particular, results like that in Theorem 1 below make it relatively easy to identify probability distributions in the credal set that “best approximate” the IM’s possibilistic output. The best approximation is relatively simple mathematically but might be difficult to compute; fortunately, further approximations can be made. Such second-level approximations can take various forms but, inspired by the asymptotic results in Martin and Williams, (2025), here I’ll use a family of Gaussian approximations as proposed in Cella and Martin, (2025). These individual Gaussian approximations on their own are not fully satisfactory, as Cella and Martin acknowledge, because they only ensure accuracy in a limited sense. If the individual approximations could be appropriately stitched together, then that would give an overall accurate approximation—but Cella and Martin weren’t able to provide a suitable stitching procedure. Here, using Theorem 1, I propose a stitching procedure that achieves the relevant statistical and computational desiderata, thereby closing the critical gap left by Cella and Martin and facilitating a principled probabilistic approximation to the IM’s output. I also show how this probabilistic approximation can be transformed back into a possibility measure that is both easy to compute and accurately approximates the target IM output.
@@
The remainder of this paper is organized as follows. Section 2 provides the necessary background on possibilistic IMs, highlighting the existing computational bottlenecks. The credal set characterization and proposed Monte Carlo strategy are presented in Section 3. Three illustrative examples, including logistic regression and a semiparametric problem involving censored data, are presented in Section 4. These are cases where direct-but-naive IM computations are expensive but not out of reach, so the accuracy of the proposed approximation can be assessed. Section 5 shows empirically that the proposed IM approximation performs well compared to standard Bayesian and likelihood-based solutions. A higher-dimensional logistic regression example is considered in Section 6, where the proposed IM approximation can, in a matter of seconds, easily solve a problem that was practically out of reach by previous efforts in the IM literature. The insights and results here in the present paper make up an important first step in a series of developments needed to create an IM toolbox for practitioners. The concluding remarks in Section 7 describe, among other things, my vision of where these developments can and need to go. Relevant technical details are given in the appendix/supplement.
@@ 2 Background
$R(z,\theta)=\frac{L_{z}(\theta)}{\sup_{\vartheta}L_{z}(\vartheta)},\quad z\in\mathbb{Z},\quad\theta\in\mathbb{T}.$
I’ll implicitly assume that the denominator is finite for almost all $z$. As is typical in the literature, I will also assume that prior information about $\Theta$ is vacuous.
The relative likelihood itself defines a data-dependent possibility contour, i.e., a non-negative function such that $\sup_{\theta}R(z,\theta)=1$ for each $z$; see Appendix A. This contour, in turn, determines a possibility measure that can be used for uncertainty quantification about $\Theta$, given $Z=z$, which has been extensively studied in the literature (e.g., Shafer, 1982; Wasserman, 1990a ; Denœux, 2006, 2014). In particular, these references suggest assigning data-driven possibility values to hypotheses $H$ about $\Theta$ via the rule
@@
$\Pi_{z}(H)=\sup_{\theta\in H}\pi_{z}(\theta),\quad H\subseteq\mathbb{T}.$ (2)
It won’t be needed in the present paper, but there’s a corresponding necessity measure, or lower probability, defined via conjugacy: $\amalg_{z}(H):=1-\Pi_{z}(H^{c})$. An essential feature of this IM construction is its so-called validity property:
@@
$\sup_{\theta\in H}\mathsf{P}_{\theta}\bigl{\{}\Pi_{Z}(H)\leq\alpha\bigr{\}}\leq\alpha,\quad\text{all $\alpha\in[0,1]$, all $H\subseteq\mathbb{T}$}.$ (5)
In words, a valid IM assigns possibility $\leq\alpha$ to true hypotheses at rate $\leq\alpha$ as a function of data $Z$. This gives the IM its “inferential weight”—(5) implies that $\Pi_{z}(H)$ is not expected to be small when $H$ is true, so one is inclined to doubt the truthfulness of a hypothesis $H$ if $\Pi_{z}(H)$ is small. Third, the above property ensures that the possibilistic IM is safe from false confidence (Balch et al., 2019; Martin, 2019; Martin, 2024b ), unlike all default-prior Bayes and fiducial solutions. An even stronger, uniform version of (5) holds, which offers opportunities for the data analyst to do more with the IM than test pre-specified hypotheses (Cella and Martin, 2023). See Martin, 2022b ; Martin, 2023b ; Martin, 2023a ; Martin, (2025) for further details about IMs’ properties, including connections to Bayesian/fiducial inference.
In a Bayesian analysis, inference is based on summaries of the posterior distribution, e.g., probabilities of relevant hypotheses, expectations of loss/utility functions, etc. All of these summaries boil down to integration involving the posterior density function. Virtually the same statement holds for the possibilistic IM: the lower–upper probability pairs for relevant hypotheses, lower–upper expectations of loss/utility functions (Martin, 2021), etc., involve optimization of the IM’s possibility contour $\pi_{z}$. More precisely, the calculus underlying possibilistic reasoning is Choquet integration (see, e.g., Troffaes and de Cooman, 2014, Appendix C). The specific details (see Appendix A) aren’t directly relevant to the developments here, so it suffices to say Choquet integration relies heavily on optimization. This observation has two important consequences: first, from the foundational point of view, compared to Bayes and other probabilistic frameworks, it’s precisely the IM’s optimization-driven Choquet integration that ensures validity is achieved; second, from a practical point of view, if reliable possibilistic inference requires optimization of $\pi_{z}$, then having efficient strategies for evaluating $\pi_{z}$ is imperative.
@@
where $Z_{m,\theta}$ are independent copies of the data $Z$, specifically drawn from $\mathsf{P}_{\theta}$, for $m=1,\ldots,M$. The above computation is feasible at one or a few different $\theta$ values, but applications often require this to be carried out over a fine grid covering the relevant portion of the parameter space. For example, identifying the confidence set in (4) requires finding which $\theta$’s satisfy $\pi_{z}(\theta)\geq\alpha$, and a naive approach is to compute the contour over a huge grid and then keep those that (approximately) meet the aforementioned condition. This amounts to lots of wasted and expensive computations. More generally, the relevant summaries of the IM output are Choquet integrals involving optimization, and doing so numerically requires many contour function evaluations. This is a serious bottleneck, so new and more efficient computational strategies are desperately needed.
@@ 3 Monte Carlo methods for IMs
3.1 Probabilistic approximation to a possibilistic IM
Given a possibilistic IM $z\mapsto(\amalg_{z},\Pi_{z})$ with contour $\pi_{z}$, I’ll refer to the level sets $C_{\alpha}(z)=\{\theta\in\mathbb{T}:\pi_{z}(\theta)\geq\alpha\}$ defined in (4) as $\alpha$-cuts. According to the theory reviewed in Section 2, the $\alpha$-cut $C_{\alpha}(z)$ is a nominal $100(1-\alpha)$confidence region for $\Theta$. Aside from the practical utility of confidence regions, there’s much more information in the IM’s $\alpha$-cuts. Towards this, the credal set $\mathscr{C}(\Pi_{z})$ corresponding to the IM’s $\Pi_{z}$ is the set of (data-dependent) probability measures that it dominates, i.e.,
$\mathscr{C}(\Pi_{z})=\{\mathsf{Q}_{z}\in\text{probs}(\mathbb{T}):\text{$\mathsf{Q}_{z}(H)\leq\Pi_{z}(H)$ for all measurable $H$}\},$
where $\text{probs}(\mathbb{T})$ is the set of probability measures on $\mathbb{T}$. Then the claim is that the $\alpha$-cuts determine the IM’s credal set. Indeed, there’s the following well-known characterization (e.g., Couso et al., 2001; Destercke and Dubois, 2014) of the credal set $\mathscr{C}(\Pi_{z})$:
$\mathsf{Q}_{z}\in\mathscr{C}(\Pi_{z})\iff\mathsf{Q}_{z}\{C_{\alpha}(z)\}\geq 1-\alpha\quad\text{for all $\alpha\in[0,1]$}.$ (7)
That is, a probability distribution $\mathsf{Q}_{z}$ belongs to the IM’s credal set if and only if it’s a confidence distribution, i.e., it assigns probability $\geq 1-\alpha$ to each of the IM’s $\alpha$-cuts. More about this notion of “confidence distributions” and when Bayesian/fiducial distributions achieve equality in (7) can be found in Martin, 2023a ; Martin, (2025). At a soon-to-be-demonstrated practical level, elements of the IM’s credal set have a relatively simple characterization, as the following theorem demonstrates. The result is presented in a more general, not-IM-specific context because this simplifies the notation.
@@ 3.1 Probabilistic approximation to a possibilistic IM
Let $\Pi$ be a possibility measure on a space $\mathbb{T}$, with corresponding contour $\pi$ and $\alpha$-cuts $C_{\alpha}=\{\theta\in\mathbb{T}:\pi(\theta)\geq\alpha\}$, $\alpha\in(0,1)$. Then $\mathsf{Q}\in\mathscr{C}(\Pi)$ if and only if
@@
See Appendix B. ∎
I would be surprised if the characterization in Theorem 1 is genuinely new, but I’ve not found a reference for this result exactly. A similar result is presented in Wasserman, 1990b (, Theorem 2.1) characterizing the credal set of a belief function determined by a given distribution and set-valued mapping. Hose, (2022, Eq. 2.44) also gives a special case of the right-hand side of (8) but makes no claims that his version is a complete characterization of $\mathscr{C}(\Pi)$; what I specifically propose to do in (10) below closely matches the Hose’s formula. As an aside, choosing the kernel $\mathsf{K}^{\beta}$ to be a uniform distribution supported on $C_{\beta}$ is an idea that commonly appears in the literature, e.g., the pignistic probability (Smets and Kennes, 1994) or the Shapley value (Shapley, 1953) of a game.
The above theorem describes the contents of the credal set $\mathscr{C}(\Pi)$ corresponding to a possibility measure $\Pi$. A relevant follow-up question is if there exists an element in $\mathscr{C}(\Pi)$ that’s “maximally consistent” with $\Pi$, i.e., a distribution $\mathsf{Q}^{\text{inn}}$ such that
@@
Such a probability distribution will be called an inner probabilistic approximation of $\Pi$. The following corollary characterizes the inner probabilistic approximations.
Corollary 1.
The equality in (9) is achieved if and only if $\mathsf{K}^{\beta}(C_{\alpha})=0$ for all pairs $(\alpha,\beta)$ with $\beta<\alpha$—that is, if $\mathsf{K}^{\beta}$ is fully supported on $\partial C_{\beta}$—and $\mathsf{M}={\sf Unif}(0,1)$.
Proof.
See the end of the proof of Theorem 1. ∎
Note that $\mathsf{Q}^{\text{inn}}$ in (9) may not exist, but this is rare in applications. Indeed, if the contour $\pi$ doesn’t span the entire range $[0,1]$, e.g., is bounded away from 0, then (9) can’t be satisfied: some of the $\alpha$-cuts $C_{\alpha}$ would be equal to $\mathbb{T}$ and every probability distribution supported on $\mathbb{T}$ would assign probability $1>1-\alpha$ to these sets. This manifests in the conditions of Corollary 1 because, in such a case, there’s no “boundary” on which to support the kernel. Even in these unfortunate cases, a probabilistic approximation can be made, but it would assign probability exceeding $1-\alpha$ to some of the $\alpha$-cuts $C_{\alpha}$.
Illustration.
As a simple-but-important application, which will prove to be useful in Section 3.3 below, consider the Gaussian possibility measure $\Pi$, with corresponding mean vector $m$ and covariance matrix $V$, whose contour is given by
@@
where $F_{d}$ is the ${\sf ChiSq}(d)$ distribution function. The corresponding $\alpha$-cuts are ellipsoids